# A Nuclear Interferometer for Ultra-Light Dark Matter Detection

### v2

@@ Abstract
We propose the nuclear interferometer - a single-photon interferometry experiment based upon the thorium-229 nuclear clock transition - as a novel detector for ultra-light dark matter (ULDM). Thanks to the enhanced sensitivity of this transition to the variation of fundamental constants, we find that possible realisations of such an experiment deploying either single ions or clouds of atoms have the potential to complement advanced very-long-baseline terrestrial clock atom interferometers in the search for ultra-light dark matter with scalar couplings to photons in the future. Nuclear interferometry may also offer an unparalleled window to new physics coupling to the QCD sector via quarks or gluons, with a discovery reach that could enhance existing and proposed experiments over a range of frequencies in the direction of well-motivated parameter space.
@@ I Introduction
ULDM could conceivably couple to the SM through a number of different ‘portals’, each of which has a distinct set of phenomenological consequences and in turn, experimental signatures. In this proposal we focus specifically on scalar ULDM with linear couplings to the SM. These couplings cause fundamental constants, and thus physical properties such as atomic energy levels and the distance scales which they govern, to oscillate in time at a frequency set by the DM mass, see e.g. Refs. [28, 29, 30, 31, 24]. This effect can be searched for in a variety of different ways including accelerometers [32], optical cavities [33], mechanical resonators [34, 35] and by comparing the frequency of two different atomic (see e.g. Refs. [29, 31, 36, 37]) or molecular [38] clocks. More recently, atom interferometers have been identified to offer the potential to probe ULDM-induced oscillations of atomic transition frequencies in a complementary region of parameter space to the aforementioned tests [39, 40]. In its most basic form, an atom interferometer constitutes an experiment which measures the phase difference between two spatially delocalised quantum superpositions of clouds of atoms [41]. Such experiments can be thought of analogously to optical interferometers but with the traditional beam-splitters and mirror elements realised via a sequence of controlled duration laser pulses which act to divide, reflect and recombine the atomic wave-packets by driving transitions between internal energy states. The presence of ULDM can be diagnosed by the existence of an oscillating phase difference between a pair of spatially separated single-photon [^1] Here, ‘single-photon’ refers specifically to a class of interferometers in which the desired transition is driven by a single laser frequency. This contrasts with interferometers based on Raman transitions which involve two distinct optical fields (see e.g. Ref. [42]). atom interferometers which are resonantly interrogated with common laser pulses [40]. Such multi-atom interferometer configurations are termed single-photon atom gradiometers and have the key advantage in that noise arising from fluctuations and jitters in the laser phase is common to both instruments and thus cancels in the differential measurement [43]. In order to achieve sensitivity to ULDM, such proposals look to both long ($\mathcal{O}$(km)) baselines and large-momentum-transfer (LMT) atom optics [^2] Such techniques (see e.g. Refs. [44, 45, 46]) effect an increase of the physical baseline by employing an increased number of laser pulses to augment the momentum transfer between the two arms of each atom interferometer and in turn the phase difference accumulated at the end of the sequence. [47] (see Sec. II.2) to increase the recorded phase difference. To sustain these practices it is essential for the pair of internal states on which the atom interferometer is based to be sufficiently long-lived such that spontaneous emission, which would otherwise exponentially degrade the sensitivity due to loss of coherence, is not significant. [^3] For a discussion of some of the potential implications of spontaneous decay on long-baseline atom interferometry in space, see Refs. [48, 49]. For the very long baselines and high number of sequential pulses demanded by ULDM searches (see e.g. [50, 51]), ultra-narrow transitions such as those deployed in optical lattice clocks are therefore typically taken advantage of. For this reason such atom interferometer realisations have earned the title of ‘large-momentum-transfer clock atom interferometry’ [52].
@@
In tandem, the discovery of an excited isomeric state in 229Th  [56, 57], has led to proposals for a ‘nuclear clock’ [58, 59, 60] based on transitions between the ground and the 229mTh excited isomer state. Arising due to accidental cancellations between contributions from strong and electromagnetic interactions [61, 62, 63], the small (8.3 eV [64]) gap in energy between the ground and excited states lies in the optical range [65] and offers an enhanced sensitivity to the variation of fundamental constants at the level of several orders of magnitude [61, 66, 67, 68] over typical optical clock transitions in atoms, ions and highly charged ions [69]. Frequency comparisons between nuclear and atomic optical clocks could offer a highly sensitive probe of ULDM in the near future [70, 71, 72, 29, 73], and the (still broad) lineshape already constrains the ULDM parameter space [74]. Due to rapid internal conversion, the excited isomeric state in electrically neutral 229Th is short-lived with a half-life of $\sim 7\,\mu$s [75]. Proposed realisations of the thorium nuclear clock for deployment in such experiments therefore look to either ionic forms [59] or to a thorium-doped lattice [76] in which internal conversion is forbidden and the isomeric state becomes long-lived with a radiative lifetime of $\sim 10^{4}$ s [70].
Whilst the nuclear clock is yet to be realised experimentally, rapid progress towards this goal has been made in recent years and especially months. The production of 229Th at ISOLDE at CERN via the decay chain of 229Fr as a new pathway [77] refined the knowledge of the nuclear transition energy from the percent to the per-mill level. Subsequently, based on the development of a new tunable VUV laser [78], the first laser excitation of the 229Th nucleus was achieved at PTB [79] in Th-doped crystals [80]. While the resolution of the transition frequency here is limited by the laser linewidth to 7 GHz [79], corresponding to a relative uncertainty of $\mathcal{O}(10^{-6})$, this precision has since further been improved to 3 GHz at UCLA [81] and, by means of a VUV frequency comb at JILA, to 300 kHz [82], suppressing the relative uncertainty to $\mathcal{O}(10^{-10})$. With such developments rapidly paving the way towards an operational nuclear clock, the time to investigate other ways in which this unique transition could be exploited to probe fundamental physics is ripe.
@@ II.1 Ultra-light scalar Dark Matter
To motivate this proposal we consider an ultra-light singlet scalar field, $\phi$, which couples linearly [^4] Note that quadratic couplings also lead to the variation of fundamental constants [83], and become relevant if linear couplings are not present, forbidden for instance by a symmetry. We leave consideration of such couplings to future work. to the SM by way of non-derivative operators. We assume that this field comprises an $\mathcal{O}(1)$ fraction [^5] An exploration of the ability of scalar couplings to generate the correct relic abundance can be found in Ref. [84]. of the local DM abundance. Working in units in which $\hbar=c=1$ and following the notational conventions [^6] In this paramaterisation, the coefficients and normalisation of each of the terms in the Lagrangian are specifically chosen such that the dimensionless couplings are all RG-invariant [85]. of Refs. [85, 29] in parameterising the dimensionless couplings relative to Newton’s gravitational constant $G_{N}$, the leading-order low energy Lagrangian, applicable to scales $\gtrsim$ 1 GeV, can be expressed in terms of a relatively minimal set of operators as
@@ II.3 Experimental Sensitivity
We now quantify the reach of a potential nuclear interferometer as an ULDM detector. In practice, over an experimental campaign of duration $T_{\textnormal{int}}$, a total of $N$ measurements of the gradiometer phase difference will be performed at a regular sampling interval $\Delta t$ =$T_{\textnormal{int}}/N$, corresponding to a constant shot repetition rate. Such measurements will contain, in addition to any possible ULDM-induced signal, contributions from background noise sources.
@@
Inserting Eq. 15 and Eq. 16 into Eq. 18, one finds that the experimental sensitivity depends on the parameters $T,$ $L$, $n$ which are set in the experimental design, in addition to the shot-noise spectral density $S_{\textnormal{n}}$ and the campaign duration $T_{\textnormal{int}}$. For a gradiometer comprising two identical interferometers, the shot-noise can be described by the white noise spectrum [93, 94, 92]
@@
Before moving on we comment on the assumption that the experimental sensitivity is shot-noise limited. Achieving shot-noise limited detection is a key design goal for any atomic sensor and requires, by definition, that all other sources of noise arising from both systematic and stochastic effects, are subdominant. Whilst this ideal has already been demonstrated experimentally for small-scale table-top atom interferometers (see e.g. Refs. [96, 97, 98]), scaling the experiments up to longer baselines brings additional considerations and challenges for noise control. Comprehensive assessments of the many sources of noise relevant to long-baseline atom interferometry (e.g. gravity gradients, laser phase noise, mechanical vibrations etc.), in addition to the various practices and techniques that may be deployed to control these, have been discussed at length in literature (see e.g. Refs. [99, 100, 54, 101]) and used to place requirements on various aspects of the experimental operation. Where not already achievable with current technologies, these stand as targets for ongoing research and development. Given that the arguments for many of these noise sources carry over straightforwardly to both of the nuclear interferometry realisations presented in this work, we do not attempt to replicate this discussion, assuming that such conditions and practices could also be successfully applied to any future nuclear interferometer experiment. Instead, in Sec. III we identify just the most significant sources of noise that are likely to impact each of the proposals outlined in this work, compare them to the shot-noise floor and discuss any resulting implications for the instrument design or experimental operation.
@@ III.1 Single Ions
We model our space-based instrument on the design considered in Refs. [104, 43, 50, 105, 106], comprising two satellites in medium Earth orbit separated by a distance of 4.4×$\times 10^{7}$ m along a single line-of-sight. We take the conservative assumption that the region over which the ions are interrogated is confined to within the interior of the satellites. This constrains the maximum wavepacket separation during the interferometric sequence, $v_{n}T$, to be less than $\sim$ 1 m [50, 104]. Whilst this inequality is linear in both $T$ and $n$, we note that due to the dependence of the ion flux and thus the shot-noise spectral density on $T$, it is preferable to deploy larger $n$. We select $n$ and $T$ to satisfy this inequality, accounting for the fact that $n$ should be even and that the total duration of the interferometer sequence should exceed the total time taken for the (sequentially fired) laser pulses to traverse the baseline, i.e. $2T>(4n-1)L/c$, where $c$ is the speed of light. Whilst the parameters used in our forecasts are representative of the reach that could be attained in such an experiment we note that further optimisation could be performed if one had a particular signal or region of parameter space in mind. Although more ambitious space-based designs in which the interrogation takes place exterior to the satellite allowing for wavepacket separations on the order of $\sim 100$ m have been proposed [99], this removes the possibility of applying magnetic shielding and would thus leave the ions subject to uncontrollable electric and magnetic fields. This would likely be a major source of gradiometer phase noise which is exacerbated by the large interrogation times associated with this set-up [^9] As will be discussed in due course, the gradiometer phase shift induced by electromagnetic field noise scales with $T^{2}$. and we therefore deem such a design unsuitable and do not consider it further.
For both the terrestrial and space-based designs we set $S_{\textnormal{n}}$ to be consistent with Eq. 19, accounting for the fact that one must have both $N_{\textnormal{s}}=1$ ions per shot and $\Delta t\geq 2T$.
@@ III.2 Neutral Atoms
where $\tau\sim 7\,\mu\textnormal{s}/\ln(2)\sim 10\,\mu$s is the (mean) lifetime of the neutral thorium isomeric excited state [75].
@@
[…]
@@
[…]
@@
The GGN spectral density for both a 100 m- and a km-scale atom interferometer with the design specifications detailed in Tab. 3 were calculated in Ref. [141] in the $1010^{-3}-$ 10 Hz frequency band by modelling the seismic field as an incoherent superposition of monochromatic plane waves propagating isotropically at the Earth’s surface. Results for both a representative low and high noise seismic environment were presented, based respectively on the empirically driven Peterson new low and high noise models [142].
@@ IV Discussion
The red shading in each of these figures delineates the parameter space which is motivated from a naturalness perspective. We derive this following Ref. [40]. The central argument here is that these scalar interactions break any shift symmetry which could keep the scalar light, such that a mass-squared contribution is expected from within the UV theory, scaling as the square of the coupling. Motivated by the scales already explored by the LHC, we conservatively take the cutoff of the IR (infrared/ low-energy) theory at which these UV-contributions are predominantly generated to be $10$ TeV and determine the parameter space in which these generic UV contributions are smaller than the scalar mass, thus avoiding the requirement of fine-tuning of different UV contributions to arrive at a given scalar mass, given a certain scalar coupling.
@@
We highlight that the deployment of the nuclear clock in an interferometry-based search for ULDM as proposed in this work is complementary to its usage in nuclear-atomic optical clock frequency comparison experiments. Indeed, as can be seen in the plots, the nuclear interferometer pushes the potential parameter space coverage closer to, and in the case of $d_{\hat{m}}$ into, more theoretically motivated regions where the couplings can be deemed technically natural.
@@
Though not studied in this work, we highlight that the nuclear interferometer is not limited to investigating ULDM with scalar couplings. In particular, the pseudo-scalar coupling of axions and axion-like particles to gluons induces oscillations in the thorium transition energy at quadratic order in the axion field [157] and would thus generate a signal in a nuclear interferometer. We leave a more detailed treatment of this scenario to future work.

### v3

@@ Abstract
We propose the nuclear interferometer - a single-photon interferometry experiment based upon the thorium-229 nuclear clock transition - as a novel detector for ultra-light dark matter. Thanks to the enhanced sensitivity of this transition to the variation of fundamental constants, we find that possible realisations of such an experiment deploying either single ions or clouds of atoms have the potential to complement advanced very-long-baseline terrestrial clock atom interferometers in the search for ultra-light dark matter with scalar couplings to photons in the future. Nuclear interferometry may also offer an unparalleled window to new physics coupling to the QCD sector via quarks or gluons, with a discovery reach that could enhance existing and proposed experiments over a range of frequencies in the direction of well-motivated parameter space.
@@ I Introduction
ULDM could conceivably couple to the SM through a number of different ‘portals’, each of which has a distinct set of phenomenological consequences and in turn, experimental signatures. In this proposal we focus specifically on scalar ULDM with linear couplings to the SM. These couplings cause fundamental constants, and thus physical properties such as atomic energy levels and the distance scales which they govern, to oscillate in time at a frequency set by the DM mass, see e.g. Refs. [28, 29, 30, 31, 24]. This effect can be searched for in a variety of different ways including accelerometers [32], optical cavities [33], mechanical resonators [34, 35] and by comparing the frequency of two different atomic (see e.g. Refs. [29, 31, 36, 37]) or molecular [38] clocks. More recently, atom interferometers have been identified to offer the potential to probe ULDM-induced oscillations of atomic transition frequencies in a complementary region of parameter space to the aforementioned tests [39, 40]. In its most basic form, an atom interferometer constitutes an experiment which measures the phase difference between two spatially delocalised quantum superpositions of clouds of atoms [41]. Such experiments can be thought of analogously to optical interferometers but with the traditional beam-splitters and mirror elements realised via a sequence of controlled duration laser pulses which act to divide, reflect and recombine the atomic wave-packets by driving transitions between internal energy states. The presence of ULDM can be diagnosed by the existence of an oscillating phase difference between a pair of spatially separated single-photon [^1] Here, ‘single-photon’ refers specifically to a class of interferometers in which the desired transition is driven by a single laser frequency. This contrasts with interferometers based on Raman transitions which involve two distinct optical fields (see e.g. Ref. [42]). atom interferometers which are resonantly interrogated with common laser pulses [40]. Such multi-atom interferometer configurations are termed single-photon atom gradiometers and have the key advantage in that noise arising from fluctuations and jitters in the laser phase or mechanical vibrations is common to both instruments and thus cancels in the differential measurement [43]. In order to achieve sensitivity to ULDM, such proposals look to both long ($\mathcal{O}$(km)) baselines and large-momentum-transfer (LMT) atom optics [^2] Such techniques (see e.g. Refs. [44, 45, 46]) effect an increase of the physical baseline by employing an increased number of laser pulses to augment the momentum transfer between the two arms of each atom interferometer and in turn the phase difference accumulated at the end of the sequence. [47] (see Sec. II.2) to increase the recorded phase difference. To sustain these practices it is essential for the pair of internal states on which the atom interferometer is based to be sufficiently long-lived such that spontaneous emission, which would otherwise exponentially degrade the sensitivity due to loss of coherence, is not significant. [^3] For a discussion of some of the potential implications of spontaneous decay on long-baseline atom interferometry in space, see Refs. [48, 49]. For the very long baselines and high number of sequential pulses demanded by ULDM searches (see e.g. [50, 51]), ultra-narrow transitions such as those deployed in optical lattice clocks are therefore typically taken advantage of. For this reason such atom interferometer realisations have earned the title of ‘large-momentum-transfer clock atom interferometry’ [52].
@@
[…]
Whilst the nuclear clock is yet to be realised experimentally, rapid progress towards this goal has been made in recent years and especially months. The production of 229Th at ISOLDE at CERN via the decay chain of 229Fr as a new pathway [77] refined the knowledge of the nuclear transition energy from the percent to the per-mill level. Subsequently, based on the development of a new tunable VUV laser [78], the first laser excitation of the 229Th nucleus was achieved at PTB [79] in Th-doped crystals [80]. While the resolution of the transition frequency here is limited by the laser linewidth to 7 GHz [79], corresponding to a relative uncertainty of $\mathcal{O}(10^{-6})$, this precision has since further been improved to 3 GHz at UCLA [81] and, by means of a VUV frequency comb at JILA, to 300 kHz [82], suppressing the relative uncertainty to $\mathcal{O}(10^{-10})$. At a temperature where the temperature shift vanishes, Ref. [83] reports a frequency stability of 280 Hz (which corresponds to a fractional frequency stability of $10^{-13}$) comparing 2 different crystals over 4 months and finds a linear dependence of the linewidth on the doping concentration of the crystal. Co-thermometry is expected to reduce the fractional uncertainty from the temperature shift below $10^{-18}$. In parallel, nuclear spectroscopy in Thorium ions of different charge provides promising prospects [84, 85]. With such developments rapidly paving the way towards an operational nuclear clock, the time to investigate other ways in which this unique transition could be exploited to probe fundamental physics is ripe.
@@ II.1 Ultra-light scalar Dark Matter
To motivate this proposal we consider an ultra-light singlet scalar field, $\phi$, which couples linearly [^5] Note that quadratic couplings also lead to the variation of fundamental constants [86, 87], becoming relevant if linear couplings are not present, forbidden for instance by a symmetry. We consider the case of the QCD axion, whose defining anomalous coupling to gluons generates an oscillatory contribution to the pion mass, and in turn the thorium transition energy, at quadratic order in the axion field, in Appendix D. to the SM by way of non-derivative operators. We assume that this field comprises an $\mathcal{O}(1)$ fraction [^6] An exploration of the ability of scalar couplings to generate the correct relic abundance via thermal misalignment can be found in Ref. [88]. of the local DM abundance. Working in units in which $\hbar=c=1$ and following the notational conventions [^7] In this paramaterisation, the coefficients and normalisation of each of the terms in the Lagrangian are specifically chosen such that the dimensionless couplings are all RG-invariant [89]. of Refs. [89, 29] in parameterising the dimensionless couplings relative to Newton’s gravitational constant $G_{N}$, the leading-order low energy Lagrangian, applicable to scales $\gtrsim$ 1 GeV, can be expressed in terms of a relatively minimal set of operators as
@@ II.3 Experimental Sensitivity
We now quantify the reach of a potential nuclear interferometer as an ULDM detector. In practice, over an experimental campaign of duration $T_{\textnormal{int}}$, a total of $N$ measurements of the gradiometer phase difference will be performed at a regular sampling interval $\Delta t$ $T_{\textnormal{int}}/N$, corresponding to a constant shot repetition rate. Such measurements will contain, in addition to any possible ULDM-induced signal, contributions from background noise sources.
@@
Inserting Eq. 15 and Eq. 16 into Eq. 18, one finds that the experimental sensitivity depends on the parameters $T,$ $L$, $n$ which are set in the experimental design, in addition to the shot-noise spectral density $S_{\textnormal{n}}$ and the campaign duration $T_{\textnormal{int}}$. For a gradiometer comprising two identical interferometers, the shot-noise can be described by the (frequency independent) white noise spectrum [97, 98, 96]
@@
Before moving on we comment on the assumption that the experimental sensitivity is shot-noise limited. Achieving shot-noise limited detection is a key design goal for any atomic sensor and requires, by definition, that all other sources of noise arising from both systematic and stochastic effects, are subdominant. The use of a gradiometer configuration (as opposed to using a single interferometer) is key in this respect, allowing common mode noise which acts identically on the upper and lower interferometers, such as laser phase noise or mechanical jitters, to be rejected with a differential measurement. Whilst shot-noise limited sensitivity has already been demonstrated experimentally in table-top atom interferometers (see e.g. Refs. [100, 101, 102]), scaling up to longer baselines brings additional considerations and challenges. In particular, coherence of the interrogated species needs to be maintained over the increased interferometry sequence duration which, for the instruments in question, is typically on the order of seconds. Comprehensive assessments of the many sources of noise relevant to long-baseline atom interferometry (e.g. gravity gradients, mechanical vibrations, black-body radiation etc.), in addition to various practices and techniques that may be deployed to control, or post-correct for these, have been discussed at length in literature (see e.g. Refs. [103, 104, 54, 105]) and used to place requirements on various aspects of the experimental operation. Where not already achievable with current technologies, these stand as targets for ongoing research and development. Given that the arguments for many of these noise sources carry over straightforwardly to both of the nuclear interferometry realisations presented in this work, we do not attempt to replicate this discussion, assuming that such conditions and practices could also be successfully applied to any future nuclear interferometer experiment. We instead focus on those aspects which are unique to the nuclear interferometer configurations presented here. In Sec. III we identify the most significant sources of noise likely to impact each of the proposals outlined in this work, compare them to the shot-noise floor and discuss any resulting implications for the instrument design or experimental operation.
@@ III.1 Single Ions
[…]
@@ III.2 Neutral Atoms
[…]
@@
The interferometer sequence depicted in Fig. 1 involves (4nn-3) $\pi$-pulses in total, all but 1 of which are intended to selectively interact with a single interferometer arm. Of these, there are 2 pulses which address an $l=1$ velocity separation, 2 of which act on an $l=n$ velocity separation and a further 4 for each $l\in\{2,...,n-1\}$. In estimating the total degradation of the sensitivity due to $\pi$-pulse inefficiencies over the course of the interferometer sequence, we assume that the laser intensity used for each $\pi$-pulse takes its maximum value such that each pulse has as short a duration and thus high as efficiency as possible. In particular we note that for $l=1$ and $l=2$ (the only pulse types required for an $n=2$ LMT sequence), the minimum pulse durations are achieved at intensity ratios $I/I_{\textnormal{sat}}$ $\sim 25$ and $\sim 105$ respectively. We additionally assume that the intensity used for the central $\pi$-pulse which is designed to simultaneously interact with both arms is sufficiently high such that its duration is negligible with respect to $\tau$. The total probability of having undergone all the desired transitions at the end of the interferometer sequence is then:
@@
Given that the reduction in $\pi$-pulse efficiency is sizeable (e.g. $P_{1}\sim 0.73$), it is preferable to deploy as few LMT kicks as possible and we thus only consider n=2n=2.141414Note that it is necessary to have $n>1$ in order that the atoms predominantly propagate in the ground state allowing for interrogation times that exceed the excited state lifetime.
@@
The GGN spectral density for both a 100 m- and a km-scale atom interferometer with the design specifications detailed in Tab. 3 were calculated in Ref. [145] in the $10^{-3}-10$ Hz frequency band by modelling the seismic field as an incoherent superposition of monochromatic plane waves propagating isotropically at the Earth’s surface. Results for both a representative low and high noise seismic environment were presented, based respectively on the empirically driven Peterson new low and high noise models [146].
@@ IV Discussion
The red shading in each of these figures delineates the parameter space which is motivated from a naturalness perspective. We derive this following Ref. [40]. The central argument here is that these scalar interactions break any shift symmetry which could keep the scalar light, such that a mass-squared contribution is expected from within the UV theory, scaling as the square of the coupling. Motivated by the scales already explored by the LHC, we conservatively take the cutoff of the IR (infrared/ low-energy) theory at which these UV-contributions are predominantly generated to be $10$ TeV and determine the parameter space in which these generic UV contributions are smaller than the scalar mass, thus avoiding the requirement of fine-tuning of different UV contributions to arrive at a given scalar mass, given a certain scalar coupling. Note, however, that this line is only intended as a rough guide. Plausible UV completions could enter at lower scales, allowing for larger natural couplings.
@@
We highlight that the deployment of the nuclear clock in an interferometry-based search for ULDM as proposed in this work is complementary to its usage in nuclear-atomic optical clock frequency comparison experiments. Indeed, as can be seen in the plots, the nuclear interferometer pushes the potential parameter space coverage closer to, and in the case of $d_{\hat{m}}$ into, more theoretically motivated regions where the couplings can be deemed technically natural. The estimated extension in reach of the nuclear interferometer over nuclear-atomic clock comparisons is, however, considerably more marginal than the relative sensitivity enhancement (to the coupling to photons) that may be achieved by traditional atom interferometers compared to table-top clock frequency comparison experiments. This follows from the afore-discussed additional challenges of using 229Th (in both atomic and ionic form) for free-fall interferometry relative to e.g. a single ion clock. Given however that substantial R&D is still required to establish the nuclear clock as a competitive frequency standard, an exploration of all potential ways to probe ULDM-induced variations of fundamental constants with 229Th is warranted at this stage. Indeed, the precise requirements on the nuclear transition (e.g. frequency stability and uncertainty etc.) are likely to be different for a nuclear interferometer (particularly in the atomic case) and a nuclear-atomic clock comparison experiment, since their sensitivity is ultimately driven by very different considerations. Given the different challenges to be overcome, it is not clear at present which experiment may be able to reach its forecasted sensitivity first. One further opportunity to probe varying fundamental constants with 229Th, albeit at higher frequencies, could be with a Mössbauer inspired set-up similar to that proposed in Ref. [161].
@@
For the purpose of illustrating the experiment, we have focused on just one specific model of ULDM. Our treatment here was not intended to be exhaustive, but to showcase the potential of the experiment and pave the way towards more complete phenomenological studies in the future. Indeed, the nuclear interferometer could be used entirely analogously to investigate any other new physics model which induces periodic variations in the thorium transition energy. Both QCD axion DM and ALP DM which couple to the SM via pseudo-scalar interactions have been shown to generate such oscillations at quadratic order in the axion field [162, 163, 164], thus making for alternative theoretical targets. The signal imprinted in a nuclear interferometer by a quadratically coupled field takes a slightly modified form compared to the linear case: to exemplify this, we make projections for (specifically) QCD axion DM in Appendix D, exploiting the quadratic coupling to hadrons which emerges below the QCD scale as a consequence of the defining axion-gluon coupling.